Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2016-04-15

Observing: Mark Magee, Ira Bar

Support Team: Weizmann

TAT: Maria Teresa Botticella, Michel Dennefeld


Weather conditions


Start of night: Thin clouds, humidity ~63%, wind speed ~5.5m/s, seeing ~1.0"

Middle of night: Cloudy, humidity ~53%, wind speed ~7.3m/s, seeing ~1.1"

End of night: Cloudy, humidity ~52%, wind speed ~6.9m/s, seeing ~1.1"


------------------------------------------------


Calibrations


Evening: PESSTO_bias, PESSTO_calib*, PESSTO_SOFI_specflat, PESSTO_SOFI_SpecArc

Morning: -


------------------------------------------------


23:11 UT: Dome opened, image analysis 

23:23 UT: Follow up: Gaia16afe, B 300sx3, VRi 200sx3, seeing ~1.0"

00:22 UT: Follow up: SN2016aqf, gr13 900s, slit 1.0"

00:42 UT: Follow up: SN2015F, Vgri 200sx9 each, seeing ~0.9-1.1"

03:30 UT: Follow up: SN2016bln, gr11+16 900s each, slit 1.0"

04:16 UT: Spectroscopic standard: EG274

04:36 UT: Follow up: SN2016blz, gr11+16 900s each, slit 1.0"

05:21 UT: Spectroscopic standard: LTT7379

05:38 UT: Classification: AT2016bmi, gr13 600s, slit 1.0"

06:03 UT: Follow up: LSQ15adm, gr11 2700s, slit 1.0"

06:53 UT: Follow up: LSQ15adm, gr16 2700s, slit 1.0"

07:53 UT: Follow up: SN2015L, gr11 2700s, slit 1.5"

08:46 UT: Follow up: SN2015L, gr16 2700s, slit 1.5"

09:44 UT: Classification: AT2016blt, gr13 900s, slit 1.0"

10:05 UT: Dome closed


*There was a problem with the first set of flats taken so we ran the OB again.


Comments